Tech Staff Meeting Notes August 10, 2015
The top priority is always repair of down stations.
NM problem stations
ET problem stations
Duty Table updated June 2014
The CERI photo and annual back to school lunch will begin at 11:30 on Friday August 21.
The next meeting to discuss SIS and the level of deagregation we want in station descriptions is scheduled for Tuesday, Aug 18 at 9:30.
The six stations in the AO network are being decommissioned and the hardware returned to the ANSS Depot. CERI offered to take over O&M for two: U40A (near Mountain Home, AR) and X40A (east of Hot Springs). Jim did a recon on U40A and its doable but will take a little work. This effort will likely take about 6 months.
2015 Preventive Maintenance
NM BB/SM & SP: 32 done & 46 to go
NM SM: 6 done & 11 to go
AG BB/SM: 5 done & 1 to go
GPS: 11 all done
Misc: 2 done & 6 to go
We stopped receiving UDP packets from GNAR. It does not appear to be a firewall. Installed a cell modem until the problem can be resolved. A fanless pc is ready to be installed to provide a smarter platform on the station end to help trouble shoot the problem. A possible workaround would be to telemeter directly to the DBB repeater at Luxora. We might want to do that anyway and avoid the need for a pc at the school.
NELE is complete.
freewave serial reset installed at 9 of 11 broadbands.
We have 11 episensors cannabalized from the K2's. We'll replace the RT147's at all 11 NM broadband stations then cull the RT147's for working sensors. So far PEBM, GNAR, GLAT, HBAR, and PENM are complete. James will install rtpd and appropriate ew modules on rhydra4 so that other episensors can be tested in the test vault prior to installation.
Menlo has no immediate need for the hardware installed at EULA and we had two small events over the weekend. So we'll keep it running for the time being.
Data taken from 2.8 event near Millington: /gaia/smeagol/dat2/Working/NM/201508/20150807_212320_MAN
CPAR/TYAR- spikes interference- repeat through each other
CHNM/NNAR - dropping out spikes- repeat through each other
HENM.HNZ and .HNE - spikes
PARM.HN* - see image
Two ISIS3 upgrades remain: ARPT, and the PLAR replacement.
RDST has a very large pole right next to the vault that will corrupt any data during strong ground motion. It needs to get moved; Dave has a couple options on the Rhodes campus (the name will change when its moved). Hopefully rather than reinstall the Guralps in the new location, we can install a Basalt instead.
GILT has no GPS in the status stream. Timing is likely bad.
The win7 replacement for mudworm was installed last week. It went well other than the year being wrong on startup that caused the data times to be stamped incorrectly.
Initial test of DBB through Wilder went well. But there's too much torque in the tower and we lose connection in the wind. Jim got permission to add a stiffener to the tower.
Jim is going to test a 950MHz link between Wilder and SFTN. If that works it will be more robust than a 2.4GHz link. This test failed.
We will install a 6 ft dish at SFTN and a 4ft dish at Wilder. That should make the link work.
James reported that the DBB link between CERI and Richardson periodically goes down for about a minute. It comes back automatically and everything seeems to catch up ok. This issue may be resolved by switching to Wilder.
The waveserver continues to have occasional very gappy periods most likely due to missed messages. Steve will investigate. Example is here.
There were daily gaps in the data of 2 to 3 hours at about 2 or 3 AM for the NMAD broadband stations. James correct the problem by resettin the nmad and lnxt aironets. Are one or both of these heading for failure?
AZTN/HDBR: Shahram submitted a proposal to TDOT to replace all the sensors with Episensors (including new cabinets, etc). We asked USGS to provide new radios so we can have a continuous buffer of all channels at AZTN. We'll then allow access to NSMP so they can extract event data.
Need to review and update the power systems in 218.
John is helping Gordana Vlahovic get NCCU back online. She has a computer that needs to be configured. Its best to use earthworm exports rather than scream to scream. Steve will help get this done.
CCNC is not going through fgtn and instead repeats through ASTN. The 950MHz shot would work so it could be converted to digital.
We're still repeating CPRT through GTTN. The plan was to put a cell modem at CPRT but John found a potential internet host in the local Library. The Library doesn't have an acceptable site for the receive antenna so John is checking with the local Fire Station as a potential internent host.
PMTN was vandalized by copper thieves. John spoke to the Sheriff who advised that once they do it once, its likely to happen again. John will look for alternative locations. We might still be able to use PMTN as a repeater/internet host. John is persuing a couple possible hosts. Mitch will "close" the dataless as of 4/13/15 and James will remove PMTN from the waveservers and heli.
SMNC has a bad horizontal. Replace L4 with S-13.
Data mostly taken from: /gaia/smeagol/dat2/Working/ET/201508/20150808_215900_MAN
I still need a larger event to be able to see problems in strong motion stations.
ASTN.HNN see image
CPRT.HNN see image
SWET.HHE-drifting see image
VHTN.EHE see image
RCGA -horrible interference
STAR Node We may have a bad intrusion alarm. Should put a better lock on the cabinet.
Need to replace the current custom reset system with pair of iboots; one at Athens and the other at Star Mountain.
Lightning at CCRT has consumed two refteks, an S-13, and several other other pieces of hardware despite having the best grounding and surge supression of any station in the network. We're going to cut our losses and downgrade to ISIS4. The reftek and episensor will find a new home, possibly at CCNC.
The unused pair of ubiquiti radios will be installed at Star Mountain. All CERI data will then go over the Ubiquiti link and the EMA can use the Aironet link to transmit the camera data in real-time continuous.
Morgantown node hardware is ready to go (needs an earthworm) and waiting on completion of the ice bridge and fiber. Delivery of ice bridget and cable rack is June 23 so we're probably looking at an August node install.
Internet, accessibility, and high noise inducing RF requires we move the
MGNC central receive. John presented several possibilities and investigated
both the PSN digitizer and the Menlo Park system. We determined that eventually
we'll need to convert all nodes to the Menlo Park system so that we can obsolesce
XP. But for now, we need to reconfigure North Carolina as in
Decommission RBNC DONE
Upgrade WSNC to broadband or S-13/Reftek and cell modem. Done.
Rotate CCNC to repeat through fgtn and add to MTTN node. Done.
Establish new node at High Peak.
Transmit MGNC to High Peak. Done
Repeat TRYN through MGNC. Too much RF, will test SMNC instead
Transmit GFM to High Peak.
Transmit TVNC to High Peak.
Transmit SMNC to High Peak. Done
GFM needs to be upgraded to ISIS4. If we can get a helicopter drop for equipment and supplies then we'll also upgrade to 3-c S-13. Should also get the helicopter to remove old hardware at the same time.
CPRT is waiting on county for permission to use Library internet; will probably be ok.
GMG needs a new vertical S-13.
Jim provided an update to the spreadsheet detailing the antenna upgrades and ethernet isolation progress at each station. All but one is complete (CJTR).
The FDL wiki is here: https://umwiki.memphis.edu/display/FDL/CERI+FDL
NSMP Station List.
The seisnet group on umdrive is at http://umdrive.memphis.edu/mwithers/seisnet You will need your uuid and password.